OpenStudy (anonymous):

2 and 1thirds + 7 halves = ?? show work

OpenStudy (anonymous):

do it roh!!!!

OpenStudy (anonymous):

dear i am thinking about it but couldnt solve sorry :(

OpenStudy (anonymous):

2+1/3+7(1/2)=12/6 +2/6 21/6=35/6